OpenStudy (anonymous):

Anyone????????????????????????????????????????????????????????????????????????????????????????????What is the solution to the system of equations? 2x - y = -3 x - 2y = -4 ( , )

OpenStudy (anonymous):

Alright lets do this together.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

Okay so first lets find x. So in the second equation all we do is add 2y to both sides. When you do that you should get: \[x = 2y - 4\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So now plug in what we got for x in the first equation. Like this: \[2(2y - 4) - y = -3\]

OpenStudy (anonymous):

Now lets solve that!

OpenStudy (anonymous):

3y - 8 = -3

OpenStudy (anonymous):

So when you do... you would get \[y = \frac{ 5 }{ 3 }\]

OpenStudy (anonymous):

Now we will plug in what we got for y in the second equation...

OpenStudy (anonymous):

So we would get: \[x - 2(\frac{ 5 }{ 3 }) = -14\]

OpenStudy (anonymous):

Now when you solve it you should get: \[x = \frac{ -32 }{ 3 }\]

OpenStudy (anonymous):

okay then what

OpenStudy (anonymous):

So our answer is: \[(\frac{ -32 }{ 3 }, \frac{ 5 }{ 3 })\]

OpenStudy (anonymous):

Thats it!

OpenStudy (anonymous):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @ashley011802

OpenStudy (anonymous):

it cant be fractions

OpenStudy (anonymous):

@Tom_Boy_Rebel it cant be fractions

OpenStudy (anonymous):

it cant be fractions

OpenStudy (anonymous):

it has to be an (x,y) no fractions allowed

OpenStudy (anonymous):

Anyone?????????????????????????????????

OpenStudy (anonymous):

HELP ME PLEASE

OpenStudy (anonymous):

I can help, I just learned this yesterday.

OpenStudy (anonymous):

please do

OpenStudy (anonymous):

Refer to the Mathematica v9 solution and plot, attached.

OpenStudy (anonymous):

omg im so sick of this no one is actually helping me

OpenStudy (anonymous):

The solution to the two linear equations given in the problem statement is where the straight lines intersect. They intersect at the following point, ( -2/3 , 5/3 ), as shown in the plot and determined in the problem solution. Don't understand why you are fixated on the idea that the solutions have to be integers.

OpenStudy (anonymous):

im doing a test they can not be fractions what do you not understand about that

OpenStudy (anonymous):

@jesus_christ_dancer

OpenStudy (anonymous):

Well if you have a friend or an instructor whom you trust, ask them to review, my attached solution for any mathematica errors and/or errors in logic.

OpenStudy (anonymous):

I do online school

OpenStudy (anonymous):

Feel free to email my solution to whom ever want for evaluation. What is the solution to the system of equations? 2x - y = -3 x - 2y = -4

OpenStudy (anonymous):

Got bail out now. Have a good day.

OpenStudy (anonymous):

the correct answer waas (-2,1)